Chapter 7 Mastery

Therrell Corporation has two divisions: Bulb Division and Seed Division. The following report is for the most recent operating period:

Bulb Division Seed Division
Sales $ 297,000 $ 198,000
Variable expenses $ 77,220 $ 41,580
Traceable fixed expenses $ 115,080 $ 81,180
Common fixed expense $ 28,320 $ 18,880

The common fixed expenses have been allocated to the divisions on the basis of sales.

Required:

a. What is the Bulb Division’s break-even in sales dollars?

b. What is the Seed Division’s break-even in sales dollars?

c. What is the company’s overall break-even in sales dollars?

(For all requirements, do not round intermediate calculations. Round your answer to the nearest whole dollar amount.)
Answer #1

Answer:

a. What is the Bulb Division’s break-even in sales dollars?

Breakeven point in sales

= Total Fixed Expenses / Contribution margin ratio

For finding out the breakeven point we need to calculate the Contribution margin ratio

Contribution margin ratio

= Contribution margin / sales

Now we will find the Bulb Division’s break-even in sales dollars as under

Bulb division

Sales

2,97,000

Less: Variable cost

77,220

Contribution margin

2,19,780

Contribution margin ratio

$219,780/$297,000*100 = 74%

Break even point in sales dollar

$115,080/74% =$155,514

______________________________________________________________________

b. What is the Seed Division’s break-even in sales dollars?

Seed division

Sales

198000

Less: Variable cost

41580

Contribution margin

1,56,420

Contribution margin ratio

$156,420/$198,000*100 = 79%

Break even point in sales dollar

$81,180/79% = $102,759

____________________________________________________________________________

c. What is the company’s overall break-even in sales dollars?

Total company

Sales

4,95,000

Less: Variable cost

1,18,800

Contribution margin

3,76,200

Contribution margin ratio

$376,200/$495,000*100 = 76%

Break even point in sales dollar

($115,080+81180+28320+18880)/76%=$320,342